OpenStudy (anonymous):

@Mathematics 6th Grade - Ratios: An airline company oversells 6 tickets for every 21 passengers. How many tickets will the airline ove0 passengers? HELP

OpenStudy (anonymous):

100

OpenStudy (anonymous):

6/21=x/350 350x6/21=100

OpenStudy (anonymous):

Explaining how to formulate the problem/solution is more helpful.

OpenStudy (anonymous):

i don't teach I answer

OpenStudy (anonymous):

tats how u get teh scholarships

OpenStudy (anonymous):

6 oversold for 21 passengers how many oversold for 350 passengers 6 for 21 x for 350 => x = 350 * 6 /21 = 2100/21 = 100